# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 4.2.1 The Short Jump
Note that for general Markov chains $p$ can be exponentially small. However, in certain cases such as single site Glauber dynamics on bounded degree graphs, $p$ is only polynomially small. Therefore, in these special cases, our main theorem implies a super quadratic speed up although the speedup term $c$ might be falling with $n$.
Properties of the Short Jump
We now show some properties of the ground state $|\psi_{b}\rangle$ that is obtained as a result of the short jump. In the next section, we will use an approximation ground state projector to bound the runtime of the long and short jumps together (as in  [DPCB23]). It is instructive however, to bound the runtime of the short jump alone and confirm that it indeed takes only polynomial time. In the following, we show that if we select $b$ such that a $\theta$-short path condition is satisfied, the time taken for the short jump is in fact $\mathcal{O}(\delta^{-1})$ where $\delta$ is the spectral gap of $P$. This analysis supports the discussion of the algorithm in Section 3. We first establish a technical condition on the change in ground state energy. An analogous bound is used in [DPCB23] to bound the runtime and while we will do the same, it is convenient to introduce it here as we establish further properties of $|\psi_{b}\rangle$.
Lemma 4.19 (Ground-state energy shift bound, adapted from Proposition 6 of [DPCB23]).
Suppose that $\gamma$-spectral density holds and the $\frac{\theta}{2}$-short-path condition holds at $b$, then
$\lvert E_{b}\rvert<1+\frac{4(\pi(E^{\ast}))^{\gamma}}{\theta}.$ (33)
The above lemma utilizes the following expression for $E_{b}$:
$E_{b}$ $=\langle\psi_{b}|H_{b}|\psi_{b}\rangle=-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-b^{2}\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle,$
where
$W_{b}$ $:=(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp})^{-1}.$
The main purpose of this lemma is for determining the cost of the long-jump, which we do in the next section. Specifically, we will want to bound the magnitude of the energy shift away from two. We provide intuition why doing so does not require a strong ($\theta$ being inverse-polynomial in $n$) short-path condition. In the above lemma, the short-path condition is used to bound the spectral norm of $W_{b}$, where a $\theta$ short-path condition implies $\lVert W_{b}\rVert_{2}\leq\theta^{-1}$. However, the actual quantity of interest is $\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle$, where $\lVert G_{\eta}|\sqrt{\pi}\rangle\rVert_{2}\leq\pi(E^{\ast})^{\gamma}$. For “hard” problems $\pi(E^{\ast})^{\gamma}$ will be exponentially small in the problem size, and so $\lVert W_{b}\rVert$ can be even exponentially large for a small ground-state energy shift. Thus, for such problems, the short-path condition we need for a good gap ($\theta$ being inverse-polynomial in $n$) is stronger than the short-path condition needed for a small-enough energy shift.
If we do have an inverse-polynomial short-path condition, then the above lemma places strong conditions on the quantum state obtained by the short jump, namely that its trace distance from the starting state is small whenever hard to solve exactly by sampling from $\pi$. Specifically, we have the following lemma
Lemma 4.20.
Suppose that the $\gamma$ spectral density and $\frac{\delta}{2}$-short path condition hold for some constant $b$. It holds that $\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\mathcal{O}\left(\delta^{-1}(\pi(E^{\ast}))^{\gamma/2}\right)$ where $\delta$ is the spectral gap of $P$. Consequently, if $\delta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right),\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\mathcal{O}^{\ast}\left((\pi(E^{\ast}))^{\gamma/2}\right)$.
Proof.
Note that the ground state energy of $H_{b}$ is $\leq-1$ since $H_{b}-(-D(P))$ is negative definite. It follows from Lemma 4.19 and the conditions of this lemma that $-1-\mathcal{O}^{\ast}(\pi(E^{\ast}))^{\gamma})\leq E_{b}\leq-1$. We observe that
$0\leq-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle-\langle\psi_{b}|H_{b}|\psi_{b}\rangle=\mathcal{O}(\delta^{-1}\pi(E^{\ast})^{\gamma}),$
$\implies$ $\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle=\mathcal{O}(\delta^{-1}\pi(E^{\ast})^{\gamma})+b\langle\psi_{b}|G_{\eta}|\psi_{b}\rangle$
$\implies$ $\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert=\mathcal{O}(\delta^{-1}(\pi(E^{\ast})^{\gamma}),$
where the last equality follows from the negative semi-definiteness of $G_{\eta}$.
We may write the state $|\psi_{b}\rangle$ as $\alpha|\sqrt{\pi}\rangle+\sqrt{1-\alpha^{2}}|\sqrt{\pi}^{\perp}\rangle$ where $|\pi^{\perp}\rangle$ is a quantum state such that $\langle\sqrt{\pi}^{\perp}|\sqrt{\pi}\rangle=0$ (We may take $\alpha$ to be real and in the interval $[0,1]$ since $|\psi_{b}\rangle,|\sqrt{\pi}\rangle$ are the ground states of stoquastic Hermitian matrices). It follows from the definition of trace distance that $\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\sqrt{1-\alpha^{2}}$. From the definition of spectral gap, and observing that ${\sqrt{\pi}}$ is the ground state of $-D(P)$, we have
$\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert$
$=$ $\lvert(1-\alpha^{2})\left(\langle\sqrt{\pi}^{\perp}|D(P)|\sqrt{\pi}^{\perp}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\right)\rvert\geq\delta(1-\alpha^{2}).$
From our two bounds on $\lvert\langle\psi_{b}|D(P)|\psi_{b}\rangle-\langle\sqrt{\pi}|D(P)|\sqrt{\pi}\rangle\rvert$, it follows that
$\lVert|\sqrt{\pi}\rangle\langle\sqrt{\pi}|-|\psi_{b}\rangle\langle\psi_{b}|\rVert_{\mathrm{Tr}}=\sqrt{1-\alpha^{2}}=\mathcal{O}\left(\delta^{-1}(\pi(E^{\ast}))^{\gamma/2}\right)=\mathcal{O}^{\ast}\left(\pi(E^{\ast}))^{\gamma/2}\right),$
which completes the proof. ∎
Since $|\sqrt{\pi}\rangle$ and $|\psi_{b}\rangle$ are pure states, we obtain a bound on the overlap between the states as an immediate consequence of the above.
Corollary 4.21.
Suppose that the $\gamma$ spectral density and $\frac{\delta}{2}$-short path condition hold for some constant $b$ with $\delta=\Omega\left(\frac{1}{\operatorname{poly}(n)}\right)$, where $\delta$ is the spectral gap of $P$. Then, $|\langle\psi_{b}|\sqrt{\pi}\rangle|=1-\mathcal{O}^{\ast}((\pi(E^{\ast})^{\gamma})$.
It is also clear from the above that if $\pi(E^{\ast})^{-1}$ is super-polynomial in $n$ (as is the case for the problems considered here), the trace distance between $|\psi\rangle$ and $|\psi_{b}\rangle$ decays super-polynomially in $n$. From the operational definition of trace distance, the difference in probability of obtaining a specific outcome when performing a fixed measurement on two states $\rho_{1},\rho_{2}$ is upper bounded by their trace distance. Since the trace distance is monotonic under discarding of identical subsystems, it also holds for all $m\in\mathbb{N}$ that $\lVert\rho_{1}^{\otimes m}-\rho_{2}^{\otimes m}\rVert_{\mathrm{Tr}}\leq\lVert\rho_{1}-\rho_{2}\rVert_{\mathrm{Tr}}$. As a consequence of the above considerations, and the monotonicity of trace distance we have the following corollary, which illustrates that the short jump by itself does not offer any advantage if used in a polynomial time approximation algorithm.
Corollary 4.22.
Suppose that the conditions of Lemma 4.20 are satisfied, and let $\mathcal{A}$ be any non-adaptive quantum or classical algorithm applied to $\operatorname{poly}(n)$ copies of an input quantum state. Each outcome of $\mathcal{A}$ that is observed with probability $p_{1}$ when the algorithm is applied to $|\psi_{b}\rangle$, must be observed with probability $p_{2}=p_{1}\pm\mathcal{O}^{\ast}((\pi(E^{\ast})^{\gamma})$.
@@ 4.2.2 The Long Jump
The long jump is the preparation of an optimal solution in $\Pi^{\ast}$ from $|\psi_{b}\rangle$, where ideally $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$. As discussed in the previous subsection, the only quantum speedup from this step is quadratic and is due to amplitude amplification. That is, the quantum short-jump plus classical sampling would cost $\mathcal{O}^{\ast}(\pi(E^{*})^{-(1-c)})$, provided that the overlap condition just mentioned is satisfied. Accordingly, a quantum short-jump applied to the discriminant of the Markov Chain in Markov Chain search could still provide a nontrivial speedup. The goal of this subsection is to determine conditions on the cost Hamiltonian in terms of the Markov Chain $\mathcal{M}$ under which $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$ holds for some constant $c$. If we combine with amplitude amplification we get the runtime stated for the generalized short-path algorithm.
In order to bound the terms in the runtime involving the ground state projector $|\psi_{b}\rangle\langle\psi_{b}|$, we follow [DPCB23] in approximating $|\psi_{b}\rangle\!\langle\psi_{b}|$ through the use of a simple $-\ell$ polynomial $\mathcal{P}_{\ell}$. This quantity is related to the runtime via
$\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert^{-1}+\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}\leq 2\lVert\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|\Pi^{\ast}\rVert_{2}^{-1}.$ (34)
Let $E_{b}$ denote the ground state energy of $H_{b}$.
Lemma 4.23 (Runtime bound by approximate projector).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain with $|\mathcal{X}|=V$. Define
$\mathcal{P}_{\ell}:=\left(\frac{H_{b}}{\lvert E_{b}\rvert}\right)^{\ell},$
and let $\omega$ be the log-Sobolev constant of $D(P)$. For either all even or all odd $\ell$ we have
$\left\langle\sqrt{\pi}\middle|\mathcal{P}_{\ell}\middle|z\right\rangle-V\left(1-\frac{\omega}{2}\right)^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z\right\rangle,$
for any assignment $z$. Note the same also holds with $\omega$ replaced by the spectral gap $\delta$.
We now recall the definition of $\alpha$-subdepolarizing from [DPCB23] but generalized to arbitrary Markov Chains.
@@
[…]
@@
As shown in the proof of Proposition 3 of [DPCB23], if $\Delta_{P}(\eta)$-stability holds, then $\alpha_{P}$-subdepolarizing is satisfied. However, the converse also holds.
Lemma 4.25 ($\Delta_{P}(\eta)$ stable $\iff\alpha_{P}$ subdepolarizing ).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The pair $(\mathcal{M},H)$ satisfies the $\Delta_{P}(\eta)$-stable if and only if it satisfies $\alpha_{P}$-subdepolarizing. They are related by the equation $\alpha_{P}=\frac{\Delta_{P}(\eta)}{\lvert E^{\ast}\rvert(1-\eta)}$.
We now briefly remark on some useful upper bounds on $\Delta_{P}(\eta)$ that we alluded to earlier. However, using too loose of an upper bound may result in the runtime analysis not indicating a speedup. Note that choice of the bound on $\Delta_{P}(\eta)$ is not actually used by Algorithm 2.
Lemma 4.26 (Upper bounds on $\Delta(\eta)$).
If $H$ has $P$ pseudo-Lipschitz norm $\lVert H\rVert_{P}$, then for $\eta\in[0,1)$,
@@
[…]
then for $\eta\in[0,1)$
@@
Our next result generalizes [DPCB23, Lemma 3], which uses $\alpha_{P}$ subdepolarizing (or equivalently $\Delta_{P}(\eta)$) to lower bound $\lvert\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z^{\ast}\rangle|$.
Lemma 4.27 (Overlap with general Markov chain).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. Given positive parameters $\eta<1$, $b<1$, $\alpha<(1-b)/2$, and integer $\ell$, suppose that $(H,g_{\eta})$ has the $\alpha$-subdepolarizing property, $3/\alpha^{2}\leq\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, and that $H_{b}$ satisfies the small-ground-energy shift condition. Define the function $F:[0,1]\mapsto[0,1]$ as $F(x):=1-x+x\ln(x)$. Let $z^{\ast}\in\{-1,+1\}^{n}$ be an optimal assignment, i.e. $H(z^{\ast})=E^{\ast}$. Then,
$\braket{\pi}{\mathcal{P}_{\ell}}{z^{\ast}}\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2}).$
@@
Then, the approximate projector can be written as
$\mathcal{P}_{\ell}=\frac{(-A-B)^{\ell}}{|E_{b}|^{\ell}}.$
By Lemma 33, if $\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, then we can take $|E_{b}|^{\ell}=\Theta(1)$ and ignore the denominator. To compute the numerator, start with
$\braket{\pi}{A}{z^{\ast}}=\braket{\pi}{z^{\ast}}=\pi(z^{\ast})^{1/2}$
$\braket{\pi}{B}{z^{\ast}}=b\pi(z^{\ast})^{1/2}.$
@@
[…]
$=b\sum_{x_{1}\dots x_{k}}\braket{\pi}{B}{x_{k}}\pi(x_{k})^{-1/2}P(x_{k-1},x_{k})\pi(x_{k-1})^{1/2}\cdots\pi(x_{1})^{-1/2}P(z^{\ast},x_{1})\pi(z^{\ast})^{1/2}$
@@
[…]
$=\sum_{x_{1},\dots}b^{\sum_{j}{x_{j}}}\dots\braket{x_{2}}{D}{x_{1}}f(H(x_{1})/E^{\ast})^{c_{1}}\braket{x_{1}}{D}{z^{\ast}}$
@@
[…]
@@
We only pick up $\pi(z^{\ast})^{1/2}$ term instead of $2^{-n/2}$ in front of the product. Therefore Propositions 15 and 16 of [DPCB23] hold. Hence, we obtain the stated result. ∎
The following uses the above results to bound the complexity of the short and long jumps. As mentioned earlier, the runtimes of the two jumps are bounded together due to the ease of analysis.
Theorem 4.28.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain, and suppose $|\mathcal{X}|=V$. If
$\ell~{}\geq~{}\max\left(\frac{3}{\alpha^{2}},\max_{z^{\ast}}4\frac{\ln(V/\sqrt{\pi(z^{\ast})})}{\omega}\right)$
and $\ell=\mathcal{O}\left(\frac{\omega}{(\pi(E^{\ast}))^{\gamma}}\right)$, then
$\lvert\langle\sqrt{\pi}|\psi_{b}\rangle\rvert^{-1}+\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}=\mathcal{O}\left([\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{b\eta}{2\ln(1/\pi(E^{\ast}))\alpha}\right)}\right).$
A direct consequence is that the overall complexity in terms of queries to block-encodings of $D(P)$ and $H$ is
$\mathcal{O}\left([\min(\textup{Gap}(D(P)),\textup{Gap}(H_{b}))]^{-1}[\pi(E^{\ast})^{-1}]^{\left(\frac{1}{2}-\frac{b\eta}{2\ln(1/\pi(E^{\ast}))\alpha}\right)}\right).$
Proof.
By Lemmas 4.23 and 4.19, for any optimal assignment $z^{\ast}$ we have
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2})-Ve^{-\omega\ell/2}.$
Consider $\ell\geq\max_{z^{\ast}}4\frac{\ln(V/\sqrt{\pi(z^{\ast}}))}{\omega}$, where the max is over optimal assignments, then
@@
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\pi(z^{\ast})^{1/2}\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)(e^{-1}-2e^{-2})-\pi(z^{\ast}).$ (35)
Note that $\exp\left(\frac{b}{\eta\alpha}F\left(1-\eta\right)\right)>1$, so the first term on the right hand side will dominate asymptotically. In fact $\frac{F(1-\eta)}{\eta}\geq\eta/2$, so we have
$\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z^{\ast}\right\rangle\geq\Omega\left(\sqrt{\pi(z^{\ast})\exp\left(\frac{b\eta}{\alpha}\right)}\right).$
This clearly gives that
@@
$\lVert\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|\Pi^{\ast}\rVert_{2}^{-1}\geq\Omega\left(\sqrt{\pi(E^{\ast})\exp\left(\frac{b\eta}{\alpha}\right)}\right).$
The result follows by using Equation (34). ∎
@@
It may not be immediately obvious that the conditions on $\ell$ are not contradicting. Here, we give intuition for why this is not the case for the typical applications of the algorithm. Note for an efficient algorithm, at the very least we will need $\omega^{-1}=\mathcal{O}(\operatorname{poly}(n))$. For hard problems, $\pi(E^{\ast})$ will be exponentially small in $n$, thus the $\mathcal{O}\left(\min_{z^{\ast}}\frac{\omega}{(\pi(z^{\ast}))^{\gamma}}\right)$ upper bound on $\ell$ will be significantly larger than the lower bound. Note $1/\alpha^{2}$ is significantly smaller than $1/\pi(E^{\ast})$ for a hard problem, for showing super-Grover runtime we will want $\alpha=\Theta(\frac{1}{\ln(1/\pi(E^{\ast}))})$ anyways. Thus it is fine to assume we have the conditions on $\ell$ stated in Theorem 4.28 in settings where the algorithm can successfully be applied.

### v2

@@ 4.2.1 The Short Jump
Note that for general Markov chains $p$ can be exponentially small. However, in certain cases such as single-site Glauber dynamics on bounded degree graphs, $p$ is only polynomially small, e.g. the antiferomagnetic $2$-spins in [CLV21, Proof of Theorem 1.1]. Therefore, in these special cases, our main theorem implies a super-quadratic speed up although the speedup term $c$ might be falling with $n$.
If the conditions of Theorem 4.11 are satisfied (potentially by using Theorems 4.14 or 4.16), then we only need to show that $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{-1}$ is exponentially smaller (say by $\pi(E^{*})^{c/2}$ for some constant $c$) than $\lVert\Pi^{\ast}|\sqrt{\pi}\rangle\rVert_{2}^{-1}$ to achieve a super-quadratic speedup over Markov chain search. This is the task of bounding the runtime of the long jump.
@@ 4.2.2 The Long Jump
The long jump is the preparation of an optimal solution in $\Pi^{\ast}$ given $|\psi_{b}\rangle$, where ideally $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$ for some constant $c>0$. As discussed in the previous subsection, the only quantum speedup from this step is quadratic and is due to amplitude amplification. That is, the quantum short-jump plus classical sampling (classical long-jump) would cost $\mathcal{O}^{\ast}(\pi(E^{*})^{-(1-c)})$, provided that the overlap condition just mentioned is satisfied. Accordingly, a quantum short-jump applied to the discriminant of the Markov chain could still provide a nontrivial speedup. The goal of this subsection is to determine conditions on the cost Hamiltonian in terms of the Markov chain $\mathcal{M}$ under which $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}\geq\pi(E^{*})^{1-c}$ holds for some constant $c$, or more generally derive a formula for a problem-size-dependent $c(n)$ in terms of problem parameters. If we combine the results of this section with the known runtime of amplitude amplification, then we get the runtime stated for the generalized short path algorithm (Equation (9)).
In order to lower bound $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}$, we follow [DPCB23] in approximating the ground-state projector $|\psi_{b}\rangle\!\langle\psi_{b}|$ through the use of a simple degree-$\ell$ polynomial. This also makes use of a technical condition that ensures the ground state energy of $H_{b}$, $E_{b}$, does not decrease too much from the ground state energy of D(P)-D(P). Unlike [DPCB23], we derive this condition without explicitly using the short path condition (Definition 4.10), simplifying some of the presentation and weakening some of the algorithm’s assumptions.
At a high-level, under various problem-specific conditions, we can bound $\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}^{2}$ by instead bounding the relative localization enhancement on optimal solutions attained by $|\psi_{b}\rangle$ over $|\sqrt{\pi}\rangle$, i.e. lower bounding
$\frac{|\langle z^{\ast}|\psi_{b}\rangle|}{|\langle z^{\ast}|\sqrt{\pi}\rangle|},$
for any optimal solution $z^{\ast}$. If the short jump from $|\sqrt{\pi}\rangle$ to $|\psi_{b}\rangle$ can be done in polynomial time, then the relative improvement in $\mathcal{O}^{\ast}$ runtime will be determined by the above ratio.
We start with the following lower bound that relates the localization of $|\psi_{b}\rangle$ on optimal solutions to $|\sqrt{\pi}\rangle$ and $H_{b}$. {restatable}[Overlap To Coupling Lower Bound]lemmalongJumpBound Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain with $|\mathcal{X}|=V$. Let $\nu$ be a lower bound on either the log-Sobolev or Poincaré constant of $P$. Suppose there exists an integer $\ell$ such that
$\frac{4\ln(V)}{\nu}\leq\ell\leq\frac{\langle\psi_{b}|\sqrt{\pi}\rangle}{b\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}.$
If $|\psi_{b}\rangle$ is the unique ground state of $H_{b}$ and the spectral gap of $H_{b}$ is lower bounded by $\frac{\nu}{2}$, then for any optimal assignment $z^{*}$:
$\langle\psi_{b}|z^{*}\rangle=\Omega\left(\langle\sqrt{\pi}|(-H_{b})^{\ell}|z^{*}\rangle\right),$
asymptotically in $V$.
We now recall the definition of $\alpha$-subdepolarizing from [DPCB23] but generalized to arbitrary Markov chains. The subdepolarizing property helps us to explicitly remove $H_{b}$ from the above bound and express things only in terms of $\pi(E^{\ast})$ and problem parameters.
@@
$\mathbb{E}_{y\underset{P}{\sim}x}\prod_{t=1}^{T}f\left(\frac{c_{t}H(y)}{E^{\ast}}\right)\geq\prod_{t=1}^{T}f\left(\frac{c_{t}(1-\alpha_{P})H(x)}{E^{\ast}}\right),$
@@
As shown in the proof of Proposition 3 of [DPCB23], if $\Delta_{P}$-stability holds, then $\alpha_{P}$-subdepolarizing is satisfied. However, the converse also holds. {restatable} $\Delta_{P}$ stable $\iff\alpha_{P}$ subdepolarizing ]lemmaequivDepol Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. The pair $(\mathcal{M},H)$ is $\Delta_{P}(\eta)$-stable if and only if it satisfies $\alpha_{P}$-subdepolarizing. They are related by the equation $\alpha_{P}=\frac{\Delta_{P}(\eta)}{\lvert E^{\ast}\rvert(1-\eta)}$.
We now briefly remark on some useful upper bounds on $\Delta_{P}(\eta)$ that we alluded to earlier. However, using too loose of an upper bound may result in the runtime analysis not indicating a speedup. Note that choice of the bound on $\Delta_{P}$ is not actually used by Algorithm 2. {restatable}[Upper bounds on $\Delta_{P}$]lemmadeltaUpper If $H$ has $P$ pseudo-Lipschitz norm $\lVert H\rVert_{P}$, then for $\eta\in(0,1)$,
@@
$\mathbb{E}_{y\underset{P}{\sim}x}[H(y)]\leq H(x)+\tilde{\Delta}_{P},$
then
@@
Our next result generalizes [DPCB23, Lemma 3], which uses $\alpha_{P}$ subdepolarizing (or equivalently $\Delta_{P}$ stability) to lower bound the coupling term $\langle\sqrt{\pi}|(-H_{b})^{\ell}|z^{*}\rangle$ in Lemma 4.2.2, removing the explicit dependence on $H_{b}$. This leads to a quantity that effectively determines how much more localized (in a relative sense) $\psi_{b}$ is on a given $z^{*}$ than $\sqrt{\pi}$. Note that obviously $\langle\sqrt{\pi}|(-H_{b})^{\ell}|z^{*}\rangle$ becomes the overlap between $|\sqrt{\pi}\rangle$ and $|z^{*}\rangle$ when $b=0$. However, one can show that this quantity is at least $\langle\sqrt{\pi}|z^{*}\rangle$ for $b>0$, but the stability condition is used to determine when it is strictly greater.
Lemma 4.19 (Enhanced Relative Localization).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. Given positive parameters $\eta<1$, $b<1$, $\alpha_{P}<(1-b)/2$, and integer $\ell>3/\alpha^{2}$, suppose that $(H,g_{\eta})$ has the $\alpha_{P}$-subdepolarizing property. Let $z^{\ast}\in\{-1,+1\}^{n}$ be an optimal assignment, i.e. $H(z^{\ast})=E^{\ast}$. Then,
$\langle\sqrt{\pi}|(-H_{b})^{\ell}|z^{*}\rangle=\Omega\left(\sqrt{\pi(z^{\ast})}e^{\frac{b\eta}{2\alpha_{P}}}\right).$
In other terms, when Lemma 4.2.2 holds:
$\frac{\langle z^{\ast}|\psi_{b}\rangle}{\langle z^{\ast}|\sqrt{\pi}\rangle}=\Omega\left(e^{\frac{b\eta}{2\alpha_{P}}}\right).$
@@
To compute the numerator $(A+B)^{\ell}$, start with
$\braket{\sqrt{\pi}|A|z^{\ast}}=\braket{\sqrt{\pi}|z^{\ast}}=\pi(z^{\ast})^{1/2}$
$\braket{\sqrt{\pi}|B|z^{\ast}}=b\pi(z^{\ast})^{1/2}.$
@@
$\braket{\sqrt{\pi}|BA^{k}|z^{\ast}}$ $=b\braket{\sqrt{\pi}|BD^{k}|z}=b\sum_{x_{1}\dots x_{k}}\braket{\sqrt{\pi}|B|x_{k}}\braket{x_{k}|D|x_{k-1}}\cdots\braket{x_{1}|D|z}$
$=b\sum_{x_{1}\dots x_{k}}\braket{\sqrt{\pi}|B|x_{k}}\pi(x_{k})^{-1/2}P(x_{k-1},x_{k})\pi(x_{k-1})^{1/2}\cdots\pi(x_{1})^{-1/2}P(z^{\ast},x_{1})\pi(z^{\ast})^{1/2}$
@@
$\braket{\sqrt{\pi}|\dots AB^{c_{3}}AB^{c_{2}}AB^{c_{1}}AB^{c_{0}}|z^{\ast}}$ $=\sum_{x_{1},\dots}\dots\braket{x_{4}|AB^{c_{3}}|x_{3}}\braket{x_{3}|AB^{c_{2}}|x_{2}}\braket{x_{2}|AB^{c_{1}}|x_{1}}\braket{x_{1}|AB^{c_{0}}|z^{\ast}}$
$=\sum_{x_{1},\dots}b^{\sum_{j}{x_{j}}}\dots\braket{x_{2}|D|x_{1}}f(H(x_{1})/E^{\ast})^{c_{1}}\braket{x_{1}|D|z^{\ast}}$
@@
$=\pi(z^{\ast})^{1/2}b^{\sum_{j=0}^{\infty}{c_{j}}}\mathbb{E}_{z^{\ast}\underset{P}{\sim}x_{1}}[(\tilde{f}(x_{1}))^{c_{1}}\mathbb{E}_{x_{1}\underset{P}{\sim}x_{2}}[(\tilde{f}(x_{2}))^{c_{2}}\cdots]]$
@@
We only pick up $\pi(z^{\ast})^{1/2}$ term instead of $2^{-n/2}$ in front of the product. Therefore Propositions 15 and 16 of [DPCB23] hold as long as $\ell>\frac{3}{\alpha^{2}}$. As observed in [DPCB23], the function $F(x)=1-x+x\ln(x)$” satisfies $\frac{F(1-\eta)}{\eta}\geq\eta/2$. Hence, we obtain the stated result. ∎
In the above proof, it is the inclusion of trajectories $|z^{*}\rangle|x_{1}\rangle|x_{2}\rangle\dots$ that walk around the landscape of the global min, defined by the filter $g_{\eta}$, that enhance the localization ($f$ does not annihilate all paths leaving $z^{*}$, unlike in unstructured search). After combining Lemmas 4.2.2 and 4.19, the proof of the following result is self-evident. This formalizes the second consequence of the previous lemma and also accounts for all $z^{\ast}\in\textup{Ran}~\Pi^{\ast}$.
Theorem 4.20 (Long Jump Cost for General Markov Chains).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain. Let $\nu$ be a lower bound on either the log-Sobolev or Poincaré constant of $P$. Given positive parameters $\eta<1$, $b<1$, $\alpha_{P}<(1-b)/2$. Suppose that $(H,g_{\eta})$ has the $\alpha_{P}$-subdepolarizing property, and that there exists an integer $\ell$ such that
@@
$\max\left(\frac{4\ln(V)}{\nu},\frac{3}{\alpha_{P}^{2}}\right)<\ell\leq\frac{\langle\psi_{b}|\sqrt{\pi}\rangle}{b\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}.$
If $|\psi_{b}\rangle$ is the unique ground state of $H_{b}$ and the spectral gap of $H_{b}$ is lower bounded by $\frac{\nu}{2}$, then
@@
$\lVert\Pi^{\ast}|\psi_{b}\rangle\rVert_{2}=\Omega\left(\sqrt{\pi(E^{\ast})}e^{\frac{b\eta}{2\alpha}}\right).$
@@
It may not be immediately obvious that the conditions on $\ell$ are not contradicting. Here, we give intuition for why this is not the case for typical applications of the algorithm. Note for an efficient algorithm, using Theorem 4.2, at the very least we will need $\omega^{-1}=\mathcal{O}(\operatorname{poly}(n))$. The overlap satisfies $|\langle\psi_{b}|\sqrt{\pi}\rangle|\geq\frac{\omega}{2}$ by Lemma 4.10. For hard problems, $\pi(E^{\ast})$ will be exponentially small in $n$ and $\gamma=\Theta(1)$. Note $\alpha_{P}^{2}$ is significantly larger than $\pi(E^{\ast})$ for a hard problem, for showing super-Grover runtime we will want $\alpha_{P}=\Theta(\frac{1}{\ln(1/\pi(E^{\ast}))})$ anyways. Thus, since $b\leq\frac{2}{3}\gamma\omega\ln(1/\pi(E^{\ast}))$, if the spectral density condition holds, then $\frac{1}{\gamma\ln(1/\pi(E^{\ast}))\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}$ will be exponentially larger than the lower bound on $\ell$. So it is fine to assume we have the conditions on $\ell$ stated in Theorem 4.20 in settings where the algorithm can successfully be applied (validating Remark 4.4).